it is estimated that 35% of households own a riding lawn mower. A sample of 14 house holds is stuided. what is the mean number of households who own a riding mower?
3.185
14
4.9
1.7847
Answer #1

solution :

given sample n=14

proportion p=35%=0.35

by using binomial condition for mean

mean=n*p =14*0.35=4.9

4.9 option C is right answer